A 2 km straight fence has posts that are set 10m on centre, meaning there is 10m from the centre of one post to the centre of the next. If the fence begins with a post and ends with a post, determine the number of posts in the entire fence
1
Expert's answer
2020-05-01T17:53:10-0400

Answer: 201

"Let's\;start\;calculation\;from\;the\;beginning\;of\;fence.\\Second\;post\;is\;10m\;far\;because\;beginning\;of\;fence\;\\and\;first\;post\;are\;in\;similar\;position.\\Third\;post\;is\;10+10\;=\;20m\;far\;from\;the\;beginning.\\So,\;nth\;post\;will\;be\;10m\times(n-1)\;far\;enough\;from\;the\;beginning.\\Let's\;calculate\;when\;10m\times(n-1)\;=\;2km.\\10m\times(n-1)\;=\;2000m\Rightarrow n-1=200\Rightarrow n=201\\So,\;there\;are\;201\;posts\;along\;the\;fence."